The sales price of a car is $12,590, which is 20% off the original price. What is the original price?
A. $14,310.40
B. $14,990.90
C. $15,290.70
D. $15,737.50
E. $16,935.80

Answer by vksarvepalli(154)

sales price of a car is $12,590, which is 20% off the original price(op)
that means op-(20/100)*op=12590
so op*(1-0.2)=12590
so original price =12590/0.8
thus original price =15737.50 $
Ans. D